OpenStudy (jazzyfa30):

Solve: |4x + 2| + 6 =0

mathslover (mathslover):

|4x+2| = - 6, is it possible?

mathslover (mathslover):

\(\mathsf{|x| \ge 0}\) It can never be less than zero. So no solution for x here.

OpenStudy (jazzyfa30):

what

OpenStudy (jazzyfa30):

@jhonyy9 can you explain it better

mathslover (mathslover):

jazzy can u tell me, what does |x| mean?

OpenStudy (jazzyfa30):

it's a variable for the number we don't know

mathslover (mathslover):

Not completely true. |x| is diff. from x. |x| means absolute value of x that is , if x is negative , then |x| will be positive. If x is positive, then |x| will be positive. Example : |-3| = 3 |-2| = 2

mathslover (mathslover):

So that means, |x| can not be negative.

mathslover (mathslover):

Getting it?

OpenStudy (jazzyfa30):

yes a little

mathslover (mathslover):

Good. so we have in the question : |4x+2| + 6 =0 that means, |4x+2| = -6 , right?

OpenStudy (jazzyfa30):

yes

mathslover (mathslover):

But as we discussed earlier, |anything| can not be negative. But here, we have |4x+2| as negative that is -6. So, it is not possible for any value of x.

mathslover (mathslover):

So I will say there is no solutions.

OpenStudy (jazzyfa30):

okay

OpenStudy (jazzyfa30):

wait so if the answer has to be positive

mathslover (mathslover):

Well actually, if it wasy |4x+2| - 6 = 0 then : |4x+2| = 6 You have further took cases and solved it.

mathslover (mathslover):

But here it is clear that, we have |4x+2| = - 6 which is not possible. So, the only point is to notice here that as it is not true for any value of x , so there will be no solutions. Lemme know where u have doubt ?

mathslover (mathslover):

Sorry have to go, but if u have further doubts, u can consult @Zarkon and @jhonyy9 .

OpenStudy (jazzyfa30):

okay @jhonyy9 aren't we suppose to remove the absolute value symbol

OpenStudy (jazzyfa30):

can some one please answer my question

OpenStudy (anonymous):

the answer is Empty set or no solution({ })because a number out of absolute value never be negative!!!!!

OpenStudy (jazzyfa30):

still confused

OpenStudy (anonymous):

[4x+2]=-6.....(after rearranging) it shows that any number is substituted in place of X it never gives negative number !!!!it is always >= 0

OpenStudy (jazzyfa30):

o now I get it thanks